OpenStudy (anonymous):

Compute the average rate of change of the function f(x)= x^2 + 2x on the interval from x = 11 to x =4

OpenStudy (anonymous):

please show all calc involved

OpenStudy (anonymous):

i dont have that option

OpenStudy (anonymous):

my options are 7, -7, 4 ,21 , 10 can you please show steps also

OpenStudy (rulnick):

f(11)=143 f(4)=24 avg rate of change FROM 11 TO 4 = (f(4) - f(11) ) / 7 = -119/7 = -17 Not one of your options. Are you sure you have the details exactly right?

OpenStudy (anonymous):

crap, x actually =1 not 11 so its x= 1 to x= 4

OpenStudy (rulnick):

much better!

OpenStudy (rulnick):

f(1)=3 answer is (24-3)/3 = 21/3 = 7

OpenStudy (anonymous):

thanks for the help

OpenStudy (rulnick):

you're welcome